# A Set of Tutorials for the LAMMPS Simulation Package [Article v1.0]

### v1

@@ 1.1 Scope
This set of tutorials consists of eight tutorials arranged in order of increasing difficulty. The novelties associated with each tutorial are briefly described below.
@@
In tutorial 6, a Monte Carlo simulation in the grand canonical ensemble is implemented to demonstrate how LAMMPS can be used to simulate an open system that exchanges particles with a reservoir.
@@ 2.1 Background knowledge
In addition, prior knowledge of the theoretical basics of molecular simulations and statistical physics is highly beneficial. Users may refer to textbooks such as Understanding Molecular Simulation by Daan Frenkel and Berend Smit [1], as well as Computer Simulation of Liquids by Michael Allen and Dominic Tildesley [2]. To better understand the fundamental concepts behind the soft matter systems simulated in these tutorials, users can also refer to Basic Concepts for Simple and Complex Liquids by Jean-Louis Barrat and Jean-Pierre Hansen [17], as well as Theory of Simple Liquids: with Applications to Soft Matter by Jean-Pierre Hansen and Ian Ranald McDonald [18]. For more resources, the SklogWiki platform provies a wide range of information information on statistical mechanics and molecular simulations [19].
@@ 2.2 Software/system requirements
The LAMMPS stable release version 29Aug2024_update2 [20] and the matching LAMMPS–GUI software version 1.6.12 are required to follow the tutorials, as they include features that were first introduced in these versions. For Linux (x86_64 CPU), macOS (BigSur or later), and Windows (10 and 11) you can download a precompiled LAMMPS package from the LAMMPS release page on GitHub [21]. Select a package with ‘GUI’ in the file name, which includes both, LAMMPS–GUI and the LAMMPS command-line executable. These precompiled packages are designed to be portable, and therefore omit support for parallel execution with MPI. Instructions for installing LAMMPS–GUI and using its most relevant features for the tutorials are provided in Appendix A.
@@ 2.3 About LAMMPS–GUI
LAMMPS–GUI is a graphical text editor, enhanced for editing LAMMPS input files and linked to the LAMMPS library, allowing it to run LAMMPS directly. The text editor functions similarly to other graphical editors, such as Notepad or Gedit, but offers the following enhancements specifically for LAMMPS:
@@
  • Monitoring of simulation progress
@@ 3.1.1 My first input
Everything that appears after a hash symbol (##\##) is a comment and ignored by LAMMPS. LAMMPS–GUI will color such comments in red. These five categories are not required in every input script and do not necessarily need to be in that exact order. For instance, the Settings and the Visualization categories could be inverted, or the Visualization category could be omitted. However, note that LAMMPS reads input files from top to bottom and processes each command immediately. Therefore, the Initialization and System definition categories must appear at the top of the input, and the Run category must appear at the bottom. Also, the specifics of some commands can change after global settings are modified, so the order of commands in the input script is important.
@@
Each LAMMPS command is accompanied by extensive online documentation that details the different options for that command. From the LAMMPS–GUI editor buffer, you can access the documentation by right-clicking on a line containing a command (e.g., units lj) and selecting «View Documentation for ‘units’». This action should prompt your web browser to open the corresponding URL for the online manual. A screenshot of this context menu is shown in Fig. 1.
@@
Figure 2: The binary mixture simulated in Tutorial 1, with the atoms of type 1 represented as small green spheres and lge atoms of type 2 as large blue spheres.
@@
The first line, region simbox (…), defines a region named simbox that is a block (i.e., a rectangular cuboid) extending from -20 to 20 units along all three spatial dimensions. The second line, create_box 2 simbox, initializes a simulation box based on the region simbox and reserves space for two types of atoms.
Note 3.2.
From this point on, any command referencing an atom type larger than 2 will trigger an error. While it is possible to allocate more atom types than needed, you must assign a mass and provide force field parameters for each atom type. Failing to do so will cause LAMMPS to terminate with an error.
The third line, create_atoms (…), generates 1500 atoms of type 1 at random positions within the simbox region. The integer 34134 is a seed for the internal random number generator, which can be changed to produce different sequences of random numbers and, consequently, different initial atom positions. The fourth line adds 100 atoms of type 2. Both create_atoms commands use the optional argument overlap 0.3, which enforces a minimum distance of 0.3 units between the randomly placed atoms. This constraint helps avoid “close contacts” between atoms, which can lead to excessively large forces and simulation instability.
@@
where $r$ is the inter-particle distance, $\epsilon_{ij}$ is the depth of the potential well that determines the interaction strength, and $\sigma_{ij}$ is the distance at which the potential energy equals zero. The indexes $i$ and $j$ refer to pairs of particle types. The fourth line, pair_coeff 1 1 1.0 1.0, specifies the Lennard-Jones coefficients for interactions between pairs of atoms of type 1: the energy parameter $\epsilon_{11}=1.0$ and the distance parameter $\sigma_{11}=1.0$. Similarly, the last line sets the Lennard-Jones coefficients for interactions between atoms of type 2, $\epsilon_{22}=0.5$, and $\sigma_{22}=3.0$.
Note 3.3.
By default, LAMMPS calculates the cross coefficients for different atom types using geometric average: $\epsilon_{ij}=\sqrt{\epsilon_{ii}\epsilon_{jj}}$, $\sigma_{ij}=\sqrt{\sigma_{ii}\sigma_{jj}}$. In the present case, $\epsilon_{12}=\sqrt{1.0\times 0.5}=0.707$, and $\sigma_{12}=\sqrt{1.0\times 3.0}=1.732$.
@@
The system is now fully parameterized, and the input is ready to compute forces. Let us complete the two remaining categories, Visualization and Run, by adding the following lines to initial.lmp:
# 4) Visualization
@@
You can now run LAMMPS. The simulation should finish quickly, and with the default settings, LAMMPS–GUI will open two windows: one displaying the console output and another with a chart. The «Output» window will display information from the executed commands, including the total energy and pressure at step 0, as specified by the thermodynamic data request. Since no actual simulation steps were performed, the «Charts» window will be empty.
@@
At this point, you can create a snapshot image of the current system using the «Image Viewer» window, which can be accessed by clicking the «Create Image» button in the «Run» menu. The image viewer works by instructing LAMMPS to render an image of the current system using its internal rendering library via the dump image command. The resulting image is then displayed, with various buttons available to adjust the view and rendering style. The image shown in Fig. 3 was created this way. This will always capture the current state of the system. Save the image for future comparisons.
@@
This tells LAMMPS to perform an energy minimization of the system. Specifically, LAMMPS will compute the forces on all atoms and then update their positions according to a selected algorithm, aiming to reduce the potential energy. By default, LAMMPS uses the conjugate gradient (CG) algorithm [30]. The simulation will stop as soon as the minimizer algorithm cannot find a way to lower the potential energy. Note that, except for trivial systems, minimization algorithms will find a local minimum rather than the global minimum.
Run the minimization and observe that LAMMPS–GUI captures the output and update the chart in real time (see Fig. 4). This run executes quickly (depending on your computer’s capabilities), but LAMMPS–GUI may fail to capture some of the thermodynamic data. In that case, use the «Preferences» dialog to reduce the data update interval and switch to single-threaded, unaccelerated execution in the «Accelerators» tab. You can repeat the run; each new attempt will start fresh, resetting the system and re-executing the script from the beginning.
@@
The potential energy, $U$, decreases from a positive value to a negative value (Figs. 4 and 5 a). Note that during energy minization, the potential energy equals the total energy of the system, $E=U$, since the kinetic energy, $K$, is zero. The initially positive potential energy is expected, as the atoms are created at random positions within the simulation box, with some in very close proximity to each other. This proximity results in a large initial potential energy due to the repulsive branch of the Lennard-Jones potential [i.e., the term in $1/r^{12}$ in Eq. (1)]. As the energy minimization progresses, the energy decreases - first rapidly - then more gradually, before plateauing at a negative value. This indicates that the atoms have moved to reasonable distances from one another.
@@
Since LAMMPS reads inputs from top to bottom, these lines will be executed after the energy minimization. Therefore, there is no need to re-initialize or re-define the system. The thermo command is called a second time to update the output frequency from 10 to 50 as soon as PART B of the simulation starts. In addition, a new thermo_style command is introduced to specify the thermodynamic information LAMMPS should print during during PART B. This adjustment is made because, during molecular dynamics, the system exhibits a non-zero temperature $T$ (and consequently a non-zero kinetic energy $K$, keyword ke), which are useful to monitor. The pe keyword represents the potential energy of the system, $E$, such that $U+K=E$.
@@
The fix nve command updates the positions and velocities of the atoms in the group all at every step. The group all is a default group that contains all atoms. The last two lines specify the value of the timestep and the number of steps for the run, respectively, for a total duration of 250 time units.
Note 3.5.
Since no other fix commands alter forces or velocities, and periodic boundary conditions are applied in all directions, the MD simulation will be performed in the microcanonical (NVE) ensemble, which maintains a constant number of particles and a fixed box volume. In this ensemble, the system does not exchange energy with anything outside the simulation box.
Run the simulation using LAMMPS. Initially, there is no equilibrium between potential and kinetic energy, as the potential energy decreases while the kinetic energy increases. After approximately 40 000 steps, the values for both kinetic and potential energy plateau, indicating that the system has reached equilibrium, with the total energy fluctuating around a certain constant value.
Now, we change the Run section to (note the smaller number of MD steps):
@@
So far, the simulation has been mostly monitored through the analysis of thermodynamic information. To better follow the evolution of the system and visualize the trajectories of the atoms, let us use the dump image command to create snapshot images during the simulation. We have already explored the «Image Viewer» window. Open it again and adjust the visualization to your liking using the available buttons. Now you can copy the commands used to create this visualization to the clipboard by either using the «Ctrl-D» keyboard shortcut or selecting «Copy dump image command» from the «File» menu. This text can be pasted into the into the Visualization section of PART B of the initial.lmp file. This may look like the following:
@@
This command tells LAMMPS to generate NetPBM format images every 100 steps. The two type keywords are for color and diameter, respectively. Run the initial.lmp using LAMMPS again, and a new window named «Slide Show» will pop up. It will show each image created by the dump image as it is created. After the simulation is finished (or stopped), the slideshow viewer allows you to animate the trajectory by cycling through the images. The window also allows you to export the animation to a movie (provided the FFMpeg program is installed) and to bulk delete those image files.
@@ 3.1.2 Improving the script
The created .data file contains all the information necessary to restart the simulation, such as the number of atoms, the box size, the masses, and the pair coefficients. This .data file also contains the final positions of the atoms within the Atoms section. The first five columns of the Atoms section correspond (from left to right) to the atom indexes (from 1 to the total number of atoms, 1150), the atom types (1 or 2 here), and the positions of the atoms $x$, $y$, $z$. The last three columns are image flags that keep track of which atoms crossed the periodic boundary. The exact format of each line in the Atoms section depends on the choice of atom_style, which determines which per-atom data is set and stored internally in LAMMPS.
Note 3.7.
@@
By visualizing the system (see Fig. 6), you may have noticed that some atoms left their original region during minimization. To start the simulation from a clean slate, with only atoms of type 2 inside the cylinder and atoms of type 1 outside the cylinder, let us delete the misplaced atoms by adding the following commands to improved.md.lmp:
@@
The compute reduce ave command is used to average the per-atom coordination number calculated by the coord/atom compute command. Compute commands are not automatically invoked; they require a “consumer” command that references the compute. In this case, the first compute is referenced by the second, and we reference the second in a thermo_style custom command (see below).
Note 3.8.
There is no need for a Settings section, as the settings are taken from the .data file.
Finally, let us complete the script by adding the following lines to improved.md.lmp:
# 4) Visualization
@@
Note 3.9.
A bulk system with periodic boundary conditions is expected to remain in place. Accordingly, when computing the temperature from the kinetic energy, we use $33N-$ 3 degrees of freedom since there is no global translation. In a drifting system, some of the kinetic energy is due to the drift, which means the system itself cools down. In extreme cases, the system can freeze while its center of mass drifts very quickly. This phenomenon is sometimes referred to as the “flying ice cube syndrome” [7].
@@
Figure 9: Snapshot of the binary mixture simulated during Tutorial 1 with atoms of type 1 colored according to their computed $1-2$ coordination number from the compute coor12, ranging from turquoise,c_coor12 = 0, to yellow, c_coor12 = 1, and red, c_coor12 = 2.

### v2

@@ 1.1 Scope
This set of tutorials consists of eight tutorials arranged in order of increasing difficulty. Although each tutorial can be read independently, information introduced in earlier tutorials is generally not repeated in detail in later ones. For this reason, we recommend that beginners follow the tutorials in order. The novelties associated with each tutorial are briefly described below.
@@
In Tutorial 6, the adsorption of a fluid in silica pores is modeled. To do so, a Monte Carlo simulation in the grand canonical ensemble is implemented to demonstrate how LAMMPS can be used to simulate an open system that exchanges particles with a reservoir.
@@ 2.1 Background knowledge
In addition, prior knowledge of the theoretical basics of molecular simulations and statistical physics is highly beneficial. Users may refer to textbooks such as Understanding Molecular Simulation by Daan Frenkel and Berend Smit [1], as well as Computer Simulation of Liquids by Michael Allen and Dominic Tildesley [2]. To better understand the fundamental concepts behind the soft matter systems simulated in these tutorials, users can also refer to Basic Concepts for Simple and Complex Liquids by Jean-Louis Barrat and Jean-Pierre Hansen [17], as well as Theory of Simple Liquids: with Applications to Soft Matter by Jean-Pierre Hansen and Ian Ranald McDonald [18]. For more resources, the SklogWiki platform provides a wide range of information on statistical mechanics and molecular simulations [19].
@@ 2.2 Software/system requirements
The LAMMPS stable release version 22Jul2025 [20] and the matching LAMMPS–GUI software version 1.7.0 are required to follow the tutorials, as they include features that were first introduced in these versions. For Linux (x86_64 CPU), macOS (BigSur or later), and Windows (10 and 11) you can download a pre-compiled LAMMPS package from the LAMMPS release page on GitHub [21]. Select a package with ‘GUI’ in the file name, which includes both, LAMMPS–GUI and the LAMMPS command-line executable. These pre-compiled packages are designed to be portable, and therefore omit support for parallel execution with MPI. Instructions for installing LAMMPS–GUI and using its most relevant features for the tutorials are provided in Appendix A.
@@ 2.3 About LAMMPS–GUI
LAMMPS–GUI is a graphical text editor, enhanced for editing LAMMPS input files and linked to the LAMMPS library, allowing it to run LAMMPS directly. The text editor is similar to other graphical editors, such as Notepad or Gedit, but offers the following enhancements for running LAMMPS:
@@
  • Editor switches working directory to that of input file
@@
  • Monitoring of simulation progress and parallelization
@@ 3.1 Tutorial 1: Lennard-Jones fluid
Figure 1: The binary mixture simulated in Tutorial 1, with the atoms of type 1 represented as small green spheres and the atoms of type 2 as large blue spheres.
@@ 3.1.1 My first input
Everything that appears after a hash symbol (#\#) is a comment and ignored by LAMMPS. LAMMPS–GUI will color such comments in red. These five categories are not required in every input script and do not necessarily need to be in that exact order. For instance, the Settings and the Monitoring categories could be inverted, or the Monitoring category could be omitted. However, note that LAMMPS reads input files from top to bottom and processes each command immediately. Therefore, the Initialization and System definition categories must appear at the top of the input, and the Run category must appear at the bottom. Also, the specifics of some commands can change after global settings are modified, so the order of commands in the input script is important.
@@
Note 3.2.
Each LAMMPS command is accompanied by extensive online documentation that lists and discusses the different options for that command. Most LAMMPS commands also have default settings that are applied if no value is explicitly specified. The defaults for each command are listed at the bottom of its documentation page. From the LAMMPS–GUI editor buffer, you can access the documentation by right-clicking on a line containing a command (e.g., units lj) and selecting «View Documentation for ‘units’». This action should prompt your web browser to open the corresponding URL for the online manual. A screenshot of this context menu is shown in Fig. 2.
@@
The first line, region simbox (…), defines a region named simbox that is a block (i.e., a rectangular cuboid) extending from -20 to 20 units along all three spatial dimensions. The second line, create_box 2 simbox, initializes a simulation box based on the region simbox and reserves space for two types of atoms. In LAMMPS, every atom is assigned an atom type property. This property selects which force field parameters (here, the Lennard-Jones parameters, $\epsilon_{ij}$ and $\sigma_{ij}$) are applied to each pair of atoms via the pair_coeff command. We discuss in Tutorial 2 how this applies to many-body pair styles, and in Tutorial 3 how this applies to Coulomb interactions.
Note 3.3.
From this point on, the number of atom types is “locked in”, and any command referencing an atom type larger than 2 will trigger an error. While it is possible to allocate more atom types than needed, you must assign a mass and provide force field parameters for each atom type. Failing to do so will cause LAMMPS to terminate with an error.
The third line, create_atoms (…), generates 1500 atoms of type 1 at random positions within the simbox region. The integer 34134 is a seed for the internal random number generator, which can be changed to produce different sequences of random numbers and, consequently, different initial atom positions. The fourth line adds 100 atoms of type 2. Both create_atoms commands use the optional argument overlap 0.3, which enforces a minimum distance of 0.3 length units between the randomly placed atoms. This constraint helps avoid “close contacts” between atoms, which can lead to excessively large forces and simulation instability. Each created atom in LAMMPS is automatically assigned a unique atom ID, which serves as a numerical identifier to distinguish individual atoms throughout the simulation. Atom IDs by default have the range from 1 to the total number of atoms, but this is not enforced. Deleting atoms, for example, causes “holes” in the list of atom IDs.
Note 3.4.
Another way to define a system in LAMMPS, besides the create_atoms commands, is to import an existing topology file containing atomic coordinates as well as, optionally, other attributes such as atomic velocities and the force field parameters using the read_data command, as shown in Tutorial 2.
@@
where $r$ is the inter-particle distance, $\epsilon_{ij}$ is the depth of the potential well that determines the interaction strength, and $\sigma_{ij}$ is the distance at which the potential energy equals zero. The indices $i$ and $j$ refer to pairs of atoms with the corresponding atom types. The fourth line, pair_coeff 1 1 1.0 1.0, specifies the Lennard-Jones coefficients for interactions between pairs of atoms that both have atom type 1: the energy parameter $\epsilon_{11}=1.0$ and the distance parameter $\sigma_{11}=1.0$. Similarly, the last line sets the Lennard-Jones coefficients for interactions between atoms of type 2, $\epsilon_{22}=0.5$, and $\sigma_{22}=3.0$.
Note 3.5.
By default, LAMMPS calculates the mixed force field coefficients for different atom types using geometric averages: $\epsilon_{ij}=\sqrt{\epsilon_{ii}\epsilon_{jj}}$, $\sigma_{ij}=\sqrt{\sigma_{ii}\sigma_{jj}}$. In the present case, $\epsilon_{12}=\sqrt{1.0\times 0.5}=0.707$, and $\sigma_{12}=\sqrt{1.0\times 3.0}=1.732$. Other rules can be selected using the pair_modify command.
@@
The system is now fully parameterized. Let us complete the two remaining categories, Monitoring and Run, by adding the following lines to initial.lmp:
# 4) Monitoring
@@
Note 3.6.
The ‘thermodynamic information’ printed by LAMMPS using thermo_style custom keywords refers to instantaneous values of the specified thermodynamic properties at each output step, not cumulative averages. However, LAMMPS also allows to reference a wide variety of custom data from compute styles, fix styles, and variables. These can be used for on-the-fly analysis, including cumulative and sliding-window averages.
You can now run LAMMPS (see subsection A.3 for details on running LAMMPS). The simulation should finish quickly, and with the default settings, LAMMPS–GUI will open two windows: one displaying the console output and another with a chart. The «Output» window will display information from the executed commands, including the total energy and pressure at step 0, as specified by the thermodynamic data request. Since no actual simulation steps were performed, the «Charts» window will be empty.
@@
At this point, you can create a snapshot image of the current system using the «Image Viewer» window, which can be accessed by clicking the «Create Image» button in the «Run» menu. The image viewer works by instructing LAMMPS to render an image of the current system using its internal rendering library via the dump image command. The resulting image is then displayed, with various buttons available to adjust the view and rendering style. The image shown in Fig. 3 was created this way. This will always capture the current state of the system. Save the image for future comparisons by clicking the «Save as» button in the «File» menu.
@@
This tells LAMMPS to perform an iterative energy minimization of the system. Specifically, LAMMPS will compute the forces on all atoms and then update their positions according to a selected algorithm using the computed forces, aiming to reduce the potential energy. By default, LAMMPS uses the conjugate gradient (CG) algorithm [30]. The simulation will stop as soon as one of the four minimizer criteria is met. LAMMPS will then report which stopping criterion was satisfied, along with selected system properties at both the initial and final steps. Note that, except for trivial systems, minimization algorithms will find a local minimum rather than the global minimum.
Run the minimization and observe that LAMMPS–GUI captures the output and update the chart in real time (see Fig. 4). This run executes quickly (depending on your computer’s capabilities) and thus LAMMPS–GUI may fail to capture some of the thermodynamic data. In that case, use the «Preferences» dialog to reduce the data update interval and switch to single-threaded, unaccelerated execution in the «Accelerators» tab. You can repeat the run; each new attempt will start fresh, resetting the system and re-executing the script from the beginning.
@@
The potential energy, $U$, decreases from a positive value to a negative value (Figs. 4 and 5 a). Note that during energy minimization, the potential energy equals the total energy of the system, $E=U$, since the kinetic energy, $K$, is zero. The initially positive potential energy is expected, as the atoms are created at random positions within the simulation box, with some in very close proximity to each other. This proximity results in a large initial potential energy due to the repulsive branch of the Lennard-Jones potential [i.e., the term $1/r^{12}$ in Eq. (1)]. As the energy minimization progresses, the energy decreases - first rapidly - then more gradually, before plateauing at a negative value. This indicates that the atoms have moved to reasonable distances from one another.
Note 3.7.
@@
Since LAMMPS reads inputs from top to bottom, these lines will be executed after the energy minimization. Therefore, there is no need to re-initialize or re-define the system. The thermo command is called a second time to update the output frequency from 10 to 50 as soon as PART B of the simulation starts. In addition, a new thermo_style command is introduced to specify the thermodynamic information LAMMPS should print during during PART B. This adjustment is made because, during molecular dynamics, the system exhibits a non-zero temperature $T$ (and consequently a non-zero kinetic energy $K$, keyword ke), which are useful to monitor. The pe keyword represents the potential energy of the system, $U$, such that $U+K=E$.
@@
The fix nve command updates the positions and velocities of the atoms in the group all at every step. More specifically, this command integrates Newton’s equations of motion using the velocity-Verlet algorithm. The group all is a default group that contains all atoms. The last two lines specify the value of the timestep and the number of steps for the run, respectively, for a total duration of 250 time units.
Note 3.8.
Since the only command affecting forces and velocities in the present script is fix nve, and periodic boundary conditions are applied in all directions, the MD simulation will be performed in the microcanonical (NVE) statistical mechanical ensemble, which maintains a constant number of particles and a fixed box volume. In this ensemble, the system does not exchange energy with anything outside the simulation box.
Run the simulation using LAMMPS. Initially, the system is not equilibrated, as the potential energy decreases while the kinetic energy increases. After approximately 40 000 steps, the values for both kinetic and potential energy plateau, indicating that the system has reached equilibrium, with the total energy fluctuating around a certain constant value.
Now, we change the second Run section to (note the smaller number of MD steps):
@@
Note 3.9.
All simulations presented in these tutorials are deliberately kept short so they can be executed on a personal computer. These runs are not intended to produce statistically meaningful results, and should not be considered suitable for publication (see for instance Ref. 32).
@@
So far, the simulation has been mostly monitored through the analysis of thermodynamic information. To better follow the evolution of the system and visualize the trajectories of the atoms, let us use the dump image command to create snapshot images during the simulation. We have already explored the «Image Viewer» window. Open it again and adjust the visualization to your liking using the available buttons. Now you can copy the commands used to create this visualization to the clipboard by either using the «Ctrl-D» keyboard shortcut or selecting «Copy dump image command» from the «File» menu. This text can be pasted into the into the Monitoring section of PART B of the initial.lmp file. This may look like the following:
@@
The ‘&\&’ characters at the end are used to extend the commands across multiple lines. These two commands tell LAMMPS to generate NetPBM format images every 100 steps. The two type keywords are for color and diameter, respectively. Run the initial.lmp using LAMMPS again, and a new window named «Slide Show» will pop up. It will show each image created by the dump image as it is created. After the simulation is finished (or stopped), the slideshow viewer allows you to animate the trajectory by cycling through the images. The window also allows you to export the animation to a movie (provided the FFMpeg program is installed) and to bulk delete those image files.
@@ 3.1.2 Improving the script
The created .data file contains all the information necessary to restart the simulation, such as the number of atoms, the box size, the masses, and the pair coefficients. This .data file also contains the final positions of the atoms, along with their IDs and types, within the Atoms section. The first five columns of the Atoms section correspond (from left to right) to the atom indexes (from 1 to the total number of atoms, 1150), the atom types (1 or 2 here), and the positions of the atoms $x$, $y$, $z$. The last three columns are image flags that keep track of which atoms crossed the periodic boundary. The exact format of each line in the Atoms section depends on the choice of atom_style, which determines which per-atom data is set and stored internally in LAMMPS.
Note 3.11.
@@
By visualizing the system (see Fig. 6), you may have noticed that some atoms left their original region during minimization. To start the simulation from a clean slate, with only atoms of type 2 inside the cylinder and atoms of type 1 outside the cylinder, let us delete the misplaced atoms by adding the following commands to the System definition section of the improved.md.lmp:
@@
Note 3.12.
The n1_in and n2_in defined above are equal-style variables, which evaluate a numerical expression using the count() function. Other LAMMPS variable styles include atom, index, file, loop, string, and vector.
@@
The compute reduce ave command is used to average the per-atom coordination number calculated by the compute coord/atom command. Compute commands do not print or output anything by themselves, nor are they automatically executed; they require a “consumer” command that references the compute. In this case, the first compute is referenced by the second, and we reference the second in a thermo_style custom command (see below).
Note 3.13.
LAMMPS compute commands can produce a wide variety of data and one can identify the category from the name of the compute style: global data (no suffix), local data (/local suffix), per-atom data (/atom suffix), per-chunk data (/chunk suffix), per-gridpoint data (/grid suffix). In the example above, the compute coord/atom produces per-atom data, which is used as input for compute reduce which returns global data. For global data three kinds of data exists: scalars (single values), vectors (one-dimensional arrays), or arrays (two-dimensional tables). When referencing results of a compute, you can use indices: for example, c_mycompute refers to the entire scalar, vector, or array, and c_mycompute[1] refers to its first element (in case of vector or array). In some cases also wildcards like “*” can be used to, for instance, refer to all elements of a vector instead of having specify all elements individually. In general, “consumer” commands (fix styles or dump styles, variables, or other compute styles) can only work with certain data types or need to have keywords set to select which data to use. You need to check the documentation of each command to ensure compatibility.
There is no need for a Settings section, as the settings are taken from the .data file. Finally, let us complete the script by adding the following lines to improved.md.lmp:
# 4) Monitoring
@@
Note 3.14.
A bulk system with periodic boundary conditions is expected to remain in place. Accordingly, when computing the temperature from the kinetic energy, we use 3N33N-3 degrees of freedom since there is no global translation. In a drifting system, some of the kinetic energy is due to the drift, which means the system itself cools down. In extreme cases, the system can freeze while its center of mass drifts very quickly. This phenomenon is sometimes referred to as the “flying ice cube syndrome” [7].
@@
Figure 9: Snapshot of the binary mixture simulated during Tutorial 1 with atoms of type 1 colored according to their computed 121-2 coordination number from the compute coor12, ranging from turquoise,c_coor12 = 0, to yellow, c_coor12 = 1, and red, c_coor12 = 2.
Note 3.15.
In contrast to the fix nve command, which integrates Newton’s equations of motion without any thermostatting, the fix nvt command adds a Nosé-Hoover thermostat to control the system temperature.